# TabPedia: Towards Comprehensive Visual Table Understanding with Concept Synergy

### v1

@@ Abstract
Tables contain factual and quantitative data accompanied by various structures and contents that pose challenges for machine comprehension. Previous methods generally design task-specific architectures and objectives for individual tasks, resulting in modal isolation and intricate workflows. In this paper, we present a novel large vision-language model, TabPedia, equipped with a concept synergy mechanism. In this mechanism, all the involved diverse visual table understanding (VTU) tasks and multi-source visual embeddings are abstracted as concepts. This unified framework allows TabPedia to seamlessly integrate VTU tasks, such as table detection, table structure recognition, table querying, and table question answering, by leveraging the capabilities of large language models (LLMs). Moreover, the concept synergy mechanism enables table perception-related and comprehension-related tasks to work in harmony, as they can effectively leverage the needed clues from the corresponding source perception embeddings. Furthermore, to better evaluate the VTU task in real-world scenarios, we establish a new and comprehensive table VQA benchmark, ComTQA, featuring approximately 9,000 QA pairs. Extensive quantitative and qualitative experiments on both table perception and comprehension tasks, conducted across various public benchmarks, validate the effectiveness of our TabPedia. The superior performance further confirms the feasibility of using LLMs for understanding visual tables when all concepts work in synergy. The benchmark ComTQA has been open-sourced at https://huggingface.co/datasets/ByteDance/ComTQA. The source code and model will be released later.
@@ 3.1 Model Architecture
Low-Resolution Vision Encoder. To keep the overall layout information, the raw image is also resized to a low-resolution one denoted as $\mathit{I}_{l}$. We choose the pre-trained CLIP visual encoder ViT-L/14 [15] to encode the low-resolution image with $224\times 224$. The output sequence $\mathit{V}_{l}$ is composed of 256 tokens, each with 512 dimension.
@@ 5.4 Ablation Studies
Impact of meditative tokens. In Tab. 10, we conduct the experiment to investigate the impact of adding meditative tokens in TabPedia. It is observed that adding meditative tokens significantly improves TabPedia’s capabilities of table perception and comprehension. We also provide a detailed analysis of the attention map of meditative tokens in Fig. D5 of Appendix. D.
@@
Table 10: Impact of vision encoders in our TabPedia.
@@
Impact of vision encoders. As shown in Tab. 10, we explore the impact of different vision encoders. In TabPedia, we propose both vision encoders to capture global and local information of the input image with different resolutions. For the high-resolution encoder, it could extract more intricate information from the text-rich image and achieve better performance than solely utilizing the low-resolution encoder. Furthermore, the low-resolution encoder plays a crucial role in furnishing comprehensive layout information, addressing the constrained receptive field of the high-resolution encoder. The results clearly indicate that the synergy between both vision encoders enhances the extraction of structural and content-related details from tables, which effectively improves perception and comprehension tasks.
@@ 6 Limitation
In this section, we discuss the limitations of our TabPedia. Firstly, since we represent the table structure with regular rectangular boxes, TabPedia is currently not capable of accurately parsing structural information for twisted or distorted tables. Secondly, all images in TQA datasets, including WTQ [82], TabFact [83] and ComTQA are dominated by tables. Therefore, TabPedia still lacks the capability to directly answer the table question with original document image. In addition, it also exhibits a deficiency in table cell recognition.
@@ Appendix B Annotation in TSR task
We illustrate the object classes utilized in TSR and TQ tasks as shown in Fig. B3. The intersection of each pair of table column and table row objects can be regarded as table grid cells. These objects construct a table’s hierarchical structure through physical overlapped rectangle boxes.

### v2

@@ Abstract
Tables contain factual and quantitative data accompanied by various structures and contents that pose challenges for machine comprehension. Previous methods generally design task-specific architectures and objectives for individual tasks, resulting in modal isolation and intricate workflows. In this paper, we present a novel large vision-language model, TabPedia, equipped with a concept synergy mechanism. In this mechanism, all the involved diverse visual table understanding (VTU) tasks and multi-source visual embeddings are abstracted as concepts. This unified framework allows TabPedia to seamlessly integrate VTU tasks, such as table detection, table structure recognition, table querying, and table question answering, by leveraging the capabilities of large language models (LLMs). Moreover, the concept synergy mechanism enables table perception-related and comprehension-related tasks to work in harmony, as they can effectively leverage the needed clues from the corresponding source perception embeddings. Furthermore, to better evaluate the VTU task in real-world scenarios, we establish a new and comprehensive table VQA benchmark, ComTQA, featuring approximately 9,000 QA pairs. Extensive quantitative and qualitative experiments on both table perception and comprehension tasks, conducted across various public benchmarks, validate the effectiveness of our TabPedia. The superior performance further confirms the feasibility of using LLMs for understanding visual tables when all concepts work in synergy. The benchmark ComTQA has been open-sourced at https://huggingface.co/datasets/ByteDance/ComTQA. The source code and model also have been released at https://github.com/zhaowc-ustc/TabPedia.
@@ 3.1 Model Architecture
Low-Resolution Vision Encoder. To keep the overall layout information, the raw image is also resized to a low-resolution one denoted as $\mathit{I}_{l}$. We choose the pre-trained CLIP visual encoder ViT-L/14 [15] to encode the low-resolution image with $224\times 224$, which has been pre-trained on 400 million image-text pairs sourced from the open-world data, thereby embedding extensive world knowledge into its pretrained weights. To preserve its generalization ability, we keep it frozen during the whole training procedure. The output sequence $\mathit{V}_{l}$ is composed of 256 tokens, each with 1024 dimension.
@@ 5.4 Ablation Studies
Necessity of Meditative Tokens. In Tab. 11, we conduct the experiment to investigate the impact of adding meditative tokens in TabPedia. It is observed that adding meditative tokens significantly improves TabPedia’s capabilities of table perception and comprehension.
What Information Matters for Meditative Tokens? We sample 100 test cases for each task and report the averaged numeric importance of high- and low-resolution vision tokens when they are attended by the meditative tokens for different tasks in the Tab. 11. Specifically, for the various VTU tasks, we calculate the averaged attention scores (across all layers and attention heads) from the LLM decoder, which indicates the extent to which the meditative tokens focus on either high- or low-resolution visual tokens. For the TSR and TQ tasks, the meditative tokens pay significantly more attention to the high-resolution visual encoder tokens. We attribute this to the fact that both tasks require more fine-grained visual information to be "deliberated" in order to construct the dense table structure. In contrast, for the TD and TQA tasks, the two visual encoders contribute almost equally to the information attended to by the meditative tokens, validating the importance of both vision encoders for these tasks.
Table 9: Contributions of different tokens.
Task
Meditative
tokens
High-res
visual tokens
Low-res
visual tokens
TD 0.65 0.16 0.19
TSR 0.64 0.12 0.24
TQ 0.71 0.11 0.19
TQA 0.56 0.18 0.25
Contributions of Different Tokens. In the Tab. 9, we calculate the averaged scores of the TabPedia-generated answers with respect to meditative tokens, high-resolution visual tokens, and low-resolution visual tokens across all the attention maps from the LLM, respectively. One can observe that the meditative tokens contribute the most information to the generation of satisfactory answers, which demonstrates that the proposed meditative tokens are indispensable and effective. We also provide a detailed analysis of the attention map of meditative tokens in Fig. D4 of Appendix. D.
@@
Table 11: Impact of different training strategies on the low-resolution vision encoder.
Task High-res visual tokens Low-res visual tokens
TD 0.49 0.51
TSR 0.71 0.29
TQ 0.73 0.27
TQA 0.51 0.49
Impact of Dual Vision Encoders. As shown in Table 13, we explore the impact of different vision encoders that capture global and local information from input images at various resolutions. The high-resolution encoder extracts intricate details from text-rich images, outperforming the low-resolution encoder, which struggles with nuanced visual representations in complex document images. Different tasks may require distinct visual cues, so dual vision encoders offer flexibility. For instance, TQA tasks need detailed table information, while TSR tasks depend on global layout. The low-resolution encoder provides comprehensive layout insights, complementing the high-resolution encoder’s limited receptive field. Our results demonstrate that combining both encoders enhances the extraction of structural and content-related details from tables, improving perception and comprehension tasks.
Frozen vs. Unfrozen Low-Resolution Vision Encoder. We further investigate different training strategies in terms of the low-resolution vision encoder. As shown in Tab. 13, it is observed that no significant performance improvement but with longer training time consumption by unfreezing it, which is in line with the conclusion in the pioneering work [97]. Besides, we suppose the encoder frozen can serve as a regularization, facilitating the extraction of layout information and alleviating potential overfitting problems, as well as more stable training. To strike the trade-off between computational consumption and performance, we thus freeze the low-resolution vision encoder during training.
Table 12: Impact of different training strategies on low-resolution vision encoder.
Low-Res Encoder PubTab1M-Det FinTabNet WTQ
Precision S-TEDS Acc
frozen 98.5 95.11 47.8
unfrozen 98.4 95.11 46.4
Table 13: Impact of dual vision encoders.
@@ 6 Limitation
In this section, we discuss the limitations of our TabPedia. Firstly, since we represent the table structure with regular rectangular boxes, TabPedia is currently not capable of accurately parsing structural information for twisted or distorted tables. Secondly, all images in TQA datasets, including WTQ [82], TabFact [83] and ComTQA are dominated by tables. Therefore, TabPedia still lacks the capability to directly answer the table question with original document image. In addition, compared to parallel decoding algorithms such as DETR [56] and Faster R-CNN [85], it consumes longer decoding time. Meantime, certain algorithmic designs such as KV cache, flash attention, and hardware improvements can effectively improve inference efficiency. We believe that with the iterative development of large model technology, the inference efficiency of TabPedia can be significantly improved.
@@ Appendix B Annotation in TSR task
We illustrate the object classes utilized in TSR and TQ tasks as shown in Fig. B3. A table generally is composed of five basic elements, i.e., column, row, spanning cell, column header and projected row header. "Row" denotes the rectangular boxes of each row’s content in the table, while "Column" denotes the rectangular boxes of each column’s content. The area where each row and each column intersect represents the table cell. Besides these both most common table elements, "Column header" refers to the area in the table that contains the data type or content for each column, usually occupying multiple rows at the top of the table. “Projected row header”, as a special row, represents the area that contains a single non-blank cell in a row. "Spanning cell" refers to a cell in a table that spans multiple rows or columns. According to these definitions, these objects have implicit relationship and construct a table’s hierarchical structure through physically overlapped rectangle boxes
@@ Appendix D More Qualitative Results
Visualization of TabPedia’s responses. As shown in Tab. D2,we introducing meditative tokens can bring promising performance across VTU tasks. We compare in detail the differences in the generated results before and after the introduction of Meditative in different VTU tasks. It is observed that introducing meditative tokens mainly improves the quality of long-form responses. Also for the perception tasks including TD and TSR, introducing meditative tokens can alleviate the meaningless or repetitive word generation. For the comprehension task, TQA, introducing meditative tokens can generate more elaborated and reasonable response. As suggested, we showcase several samples for better understanding.
Table D2: Qualitative results of TabPedia’s responses.
Image Question GT TabPedia (w/o Meditative Tokens) TabPedia
[…]